A one-year call option contract on Cheesy Poofs Co. stock sells for $1,350. In one year, the stock will be worth $67 or $88 per share. The exercise price on the call option is $80. What is the current value of the stock if the risk-free rate is 4 percent? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

C0 = $1,350/(100 shares per contract) = $13.5 per share

S0 = (SH - SL) / (SH - X)* C0 + SL/(1+R)

S0 = ($88-$67) / (88-80) ´ $13.5 + $67 / (1.04) = $35.44 + $ 64.42 = $99.86

Where, SH is higher stock price.

SL is lower stock price

X is  exercise price on the call option

C0 is value of call option as per contract

R is risk free rate.

hence current value of stock is $99.86